OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 A. x + y = y + 2 B. (y + 2) + y = 6 C. x = (x – 6) + 2 D. x + (y + 2) = 6

OpenStudy (anonymous):

I believe it is D

OpenStudy (anonymous):

@fallenangelorchid sorry but I have a lot of questions guys @pythonwizard @Alchemista @DJ3strella @AlexandervonHumboldt2 @hotdancingbabe14

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

its not d its b. you have it confused

OpenStudy (anonymous):

your substituting y+2 for x in the first eqquation

OpenStudy (karon86):

Look at the system. Notice the system literally gives you what x is just like when you solve an equation and you get x= You have x=y-2 So plug x into the other equation y+2+y=6 or (y+2)+y=6

OpenStudy (anonymous):

sowy if I do xD

OpenStudy (karon86):

SO its B

OpenStudy (anonymous):

I thought it was c at first how is it b @pythonwizard

OpenStudy (anonymous):

Yup it is B

OpenStudy (anonymous):

Karon explains it perfectly

OpenStudy (anonymous):

in these types of problems, solve for the easiest vaiable. and then plug-and-chug into the other equation. and tada, you have the answer

OpenStudy (karon86):

LOL thank you

OpenStudy (anonymous):

look at karon's reply for clarification. :-)

OpenStudy (anonymous):

is anyone here taking pre-calc at the moment?

OpenStudy (anonymous):

@thesmartone

OpenStudy (anonymous):

Sorry just Algebra 2 honors

OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x + y = 7 9x + _ y = _

OpenStudy (anonymous):

3 in the y and 21 after teh =

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

if you noticed, the bottom is just the same as the top exept everything is multiplied by 3

OpenStudy (anonymous):

so its 9x + 3y = 21 ?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

yay thanks for the tip :)

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

A collection of quarters and nickels is worth $14.80. There are 88 coins in all. How many of each coin are there? _ dimes _ nickles

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

x is the number of dimes, nickel

OpenStudy (anonymous):